Grade 12Mechanics

A position dependent force F= 7-2x+3x^2N acts on a small body of mass 2kg and displaces it from x=0 to x=5m the work done in joule is?

Raksha swami
9 Years agoGrade 12

2 Answers

Vikas TU
9 Years ago
  •  
F =7-2x+3x^2
F = dW/dx
F.dx = dW
Integrate on both sides
∫F.dx = ∫dw
W =∫05(7-2x+3x2) dx
     = [7x-2(x2)/2 + 3(x3)/3]50
     = [35 -25+125]
     =135 joule.
Kushagra Madhukar
5 Years ago
Dear student,
Please find the attached answer to your problem below.
 
F =7-2x+3x^2
F = dW/dx
F.dx = dW
Integrate on both sides
∫F.dx = ∫dw
W =∫05(7-2x+3x2) dx
     = [7x-2(x2)/2 + 3(x3)/3]50
     = [35 -25+125]
     =135 joule.
 
Hope it helps.
Thanks and regards,
Kushagra